# Implications of computer science theory for the simulation hypothesis

### v2

@@ Appendix A Appendix A: Turing machines
There are many different definitions of Turing machines that are computationally equivalent to one another, in that any computation that can be done with one type of Turing machine can be done with the other. It also means that the “scaling function” of one type of Turing machine, mapping the size of a computation to the minimal amount of resources needed to perform that computation by that type of Turing machine, is at most a polynomial function of the scaling function of any other type of Turing machine. (See for example the relation between the scaling functions of single-tape and multi-tape Turing machines [5].) The following definition will be useful for our purposes, even though it is more complicated than strictly needed:
@@ A Turing machine (TM) is a 7-tuple $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ where:
    $\rho:R\times{\mathbb{Z}}\times\Lambda^{\infty}\rightarrow R\times{\mathbb{Z}}\times\Lambda^{\infty}$ is the update function. It is required that for all triples $(r,v,T)$, that if we write $(r^{\prime},v^{\prime},T^{\prime})=\rho(r,v,T)$, then $v^{\prime}$ does not differ by more than $1$ from $v$, and the vector $T^{\prime}$ is identical to the vectors $T$ for all components with the possible exception of the component with index $v$; [^15] Technically the update function only needs to be defined on the “finitary” subset of ××${\mathbb{R}}\times{\mathbb{Z}}\times\Lambda^{\infty}$, namely, those elements of ××${\mathbb{R}}\times{\mathbb{Z}}\times\Lambda^{\infty}$ for which the tape contents has a non-blank value in only finitely many positions.
@@
Any TM $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ starts with $r=r^{\varnothing}$, the counter set to a specific initial value (e.g, 00), and with $\lambda$ consisting of a finite contiguous set of non-blank symbols, with all other symbols equal to $b$. The TM operates by iteratively applying $\rho$, if and until the computational state falls in $r^{A}$, at which time the process stops, i.e., any ID with the head in the halt state is a fixed point of $\rho$.
@@
Write the set of non-blank symbols of $\Lambda$ as $\hat{\Lambda}$. Every $\lambda$ on the tape of a TM that it might have during a computation in which it halts is a finite string of elements in $\Lambda$ delimited by an infinite string of blanks. Accordingly, wolog we often refer to the state space of the tape is $\Lambda^{*}$, with the trailing infinite string of blanks implicit. Note that $\Lambda^{*}$ is countably infinite, in contrast to $\Lambda^{\infty}$.
@@
We say that a total function $f$ from $(\Lambda\setminus\{b\})^{*}$ to itself is recursive, or (total) computable, if there is a TM with input alphabet $\Lambda$ such that for all $x\in(\Lambda\setminus\{b\})^{*}$, the TM computes $f(x)$. If $f$ is instead a partial function, then we say it is partial recursive (partial computable, resp.) if there is a TM with input alphabet $\Lambda$ that computes $f(x)$ for all $x$ for which $f(x)$ is defined, and does not halt for any other $x$. (The reader should be warned that in the literature, the term “computable” is sometimes taken to mean partial computable rather than total computable — and in some articles it is sometimes taken to mean either total computable or partial computable, depending on the context.)
@@
As mentioned, there are many variants of the definition of TMs provided above. In one particularly popular variant the single tape in Definition 7 is replaced by multiple tapes. Typically one of those tapes contains the input, one contains the TM’s output (if and) when the TM halts, and there are one or more intermediate “work tapes” that are in essence used as scratch pads. The advantage of using this more complicated variant of TMs is that it is often easier to prove theorems for such machines than for single-tape TMs. However, there is no difference in their computational power. More precisely, one can transform any single-tape TM into an equivalent multi-tape TM (i.e., one that computes the same partial function), as well as vice-versa [5, 36, 57].
@@
As is standard, I fix some set of (prefix free) encodings of all tuples of finite bit strings, $\langle$.$\rangle$ . ⟩, .,$\langle$.,$\rangle$ . , . ⟩… In particular, in general the input to a UTM is encoded as $\langle k,x\rangle$ if it is emulating TM $T^{k}$ running on input string $x$. However, sometimes for clarity of presentation I will leave the angle brackets implicit, and simply write the UTM $U$ operating on input $\langle k,x\rangle$ as $U(k,x)$. In addition, as shorthand, if $x$ is a vector whose components are all bit strings, I will write the encoded version of all of its components as $\langle x\rangle$.
@@
(where “$O(1)$” means a term that is independent of both $x$ and $y$). Indeed, in a certain technical sense, the expected value of $K(x)$ under any distribution $P(x\in\Lambda^{*})$ equals the Shannon entropy of $P$. (See [36].)
Formally speaking, the set ${\mathbb{B}}^{*}$ is a Cantor set. A convenient probability measure on this Cantor set, sometimes called the fair-coin measure, is defined so that for any binary string $x$ the set of sequences that begin with $\sigma$ has measure $2^{-|\sigma|}$. Loosely speaking, the fair-coin measure of a prefix TM $T$ is the probability distribution over the strings in $T$’s halting set generated by IID “tossing a coin” to generate those strings, in a Bernoulli process, and then normalizing. [^17] Kraft’s inequality guarantees that since the set of strings in the halting set is a prefix-free set, the sum over all its elements of their probabilities cannot exceed $1$, and so it can be normalized. However, in general that normalization constant is uncomputable, as discussed below. Also, in many contexts we can actually assign arbitrary non-zero probabilities to the strings outside the halting set, so long as the overall distribution is still normalizable. See [36]. So any string $\sigma$ in the halting set has probability $2^{-|\sigma|}/\Omega$ under the fair-coin prior, where $\Omega$ is the normalization constant for the TM in question.
The fair-coin prior provides a simple Bayesian interpretation of Kolmogorov complexity: Under that prior, the Kolmogorov complexity of any string $\sigma$ for any prefix TM $T$ is just (the log of) the maximum a posterior (MAP) probability that any string $\sigma^{\prime}$ in the halting set of $T$ was the input to $T$, conditioned on $\sigma$ being the output of that TM. (Strictly speaking, this result is only true up to an additive constant, given by the log of the normalization constant of the fair-coin prior for $T$.)
@@
I end this appendix with some terminological comments and definitions that will be useful in the main text. It is conventional when dealing with Turing machines to implicitly assume some invertible map $h(.)$ from ${\mathbb{Z}}$ into $\Lambda^{*}$. Given such a map $h(.)$, we can exploit it to implicitly assume an additional invertible map taking $\mathbb{Q}$ into $\Lambda^{*}$, e.g., by uniquely expressing any rational number as one product of primes, $a$, divided by a product of different primes, $b$; invertibly mapping those two products of primes into the single integer $2^{a}3^{b}$; and then evaluating $Rh(2^{a}3^{b})$. Using these definitions, we say that a real number $z$ is computable iff there is a recursive function $f$ mapping rational numbers to rational numbers such that for all rational-valued accuracies $\epsilon>0$, $|f(\epsilon)-z|<\epsilon$. We define computable functions from $\mathbb{Q}\rightarrow{\mathbb{R}}$ similarly.
Appendix B Appendix B: The recursion theorem and Rice’s theorem
@@
For any partial computable function $Q(x,y)$ there is a Turing machine with index $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$.
@@
For any TM of the form $M(x,y)$ taking two arguments, define $\llbracket{M(y,x)}\rrbracket_{x}$ as the index $e$ such that $T^{e}(x)=M(y,x)$ for all $x$, with $y$ fixed.
@@
Note that $S(.)$ is a total computable function. Using Eq. 37, choose an index $k$ such that
$T^{k}(t,x)=Q(S(t),y)$ (38)
(Such an index must exist since the RHS of Eq. 38 is a partial computable function.)
@@
An important special case of the theorem is where $Q(.,.)$ is a total computable function. In this case the theorem says that that there must be an $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$, and therefore is also total computable. So, we can restrict to total computable functions $Q(.,)$ and total TM’s $T^{e}$, as a special case. I call this extension the “total recursion theorem” in the main text.
The theorem has elicited some interesting commentary. For example, [44] writes that
@@
An important special case is where $G$ is restricted to a set of binary-valued partial functions. In this case Rice’s theorem concerns the decidability of sets of languages.
Appendix C Appendix C: Why $\Delta t$ is not a physical variable
@@
Ultimately, the way we are getting around these problems in the framework I’ve adopted is by having $\Delta t$ just be a parameter of the evolution function, specifying how far into the future we want to simulate the evolution of the physical universe. and not a physical variable. So for example it does not need to be reproduced as the output of the computer if the cosmological universe is to simulate its own future.) A consequence of this workaround is that we need to hard-code $n^{*}$ into the initial state of the program in a way that depends on $\Delta t$ [^19] Indeed, the way the derivation of the self-simulation lemma uses the recursion theorem can be viewed as a special case of the generalized parameter-dependent form of the recursion theorem given in [44]. The derivation of the self-simulation lemma uses that generalized form for the special case where the space of parameters is single-dimensional. So in the notation of that paper, here $m=1$.
Appendix D Appendix D: Subtleties with the model in Section 2.6
Section 2.6 presented an example of how a portion of our actual universe could implement the purely formal model of universe $V$ given in Section 2.2. In that example the computer was implemented as a UTM, replete with tapes, etc.. This might seem a particularly awkward way of modeling a portion of our cosmological universe whose dynamics is computationally universal. After all, our universe “runs in parallel”, whereas a UTM is a serial system. This means that to implement such a UTM would require copious use of energy barriers and the like, to prevent the parallel nature of Hamilton’s equations from “leaking through”.

### v3

@@ Appendix A Turing machines
There are many different definitions of Turing machines that are computationally equivalent to one another, in that any computation that can be done with one type of Turing machine can be done with the other. In fact, the “scaling function” of one type of Turing machine, mapping the size of a computation to the minimal amount of resources needed to perform that computation by that type of Turing machine, is at most a polynomial function of the scaling function of any other type of Turing machine. (See for example the relation between the scaling functions of single-tape and multi-tape Turing machines [5].) The following definition will be useful for our purposes, even though it is more complicated than strictly needed:
@@ A Turing machine (TM) is a 7-tuple $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ where:
    $\rho:R\times{\mathbb{Z}}\times\Lambda^{\infty}\rightarrow R\times{\mathbb{Z}}\times\Lambda^{\infty}$ is the update function. It is required that for all triples $(r,v,T)$, that if we write $(r^{\prime},v^{\prime},T^{\prime})=\rho(r,v,T)$, then $v^{\prime}$ does not differ by more than $1$ from $v$, and the vector $T^{\prime}$ is identical to the vectors $T$ for all components with the possible exception of the component with index vv; [^21] Technically the update function only needs to be defined on the “finitary” subset of ${\mathbb{R}}\times{\mathbb{Z}}\times\Lambda^{\infty}$, namely, those elements of ${\mathbb{R}}\times{\mathbb{Z}}\times\Lambda^{\infty}$ for which the tape contents has a non-blank value in only finitely many positions.
@@
Any TM $(R,\Lambda,b,v,r^{\varnothing},r^{A},\rho)$ starts with $r=r^{\varnothing}$, the counter set to a specific initial value (e.g, 0), and with $\lambda$ consisting of a finite contiguous set of non-blank symbols, with all other symbols equal to $b$. The TM operates by iteratively applying $\rho$, if and until the computational state falls in $r^{A}$, at which time the process stops, i.e., any ID with the head in the halt state is a fixed point of $\rho$.
@@
Write the set of non-blank symbols of $\Lambda$ as $\hat{\Lambda}$. Every $\lambda$ on the tape of a TM that it might have during a computation in which it halts is a finite string of elements in $\Lambda$ delimited by an infinite string of blanks. Accordingly, wolog we often refer to the state space of the tape as $\Lambda^{*}$, with the trailing infinite string of blanks implicit. Note that $\Lambda^{*}$ is countably infinite, in contrast to $\Lambda^{\infty}$.
@@
We say that a total function $f$ from $(\Lambda\setminus\{b\})^{*}$ to itself is recursive, or (total) computable, if there is a TM with input alphabet $\Lambda$ such that for all $x\in(\Lambda\setminus\{b\})^{*}$, the TM computes $f(x)$. If $f$ is instead a partial function, then we say it is partial recursive (partial computable) if there is a TM with input alphabet $\Lambda$ that computes $f(x)$ for all $x$ for which $f(x)$ is defined, and does not halt for any other $x$. (The reader should be warned that in the literature, the term “computable” is sometimes taken to mean partial computable rather than total computable — and in some articles it is sometimes taken to mean either total computable or partial computable, depending on the context.)
@@
As mentioned, there are many variants of the definition of TMs provided above. In one particularly popular variant the single tape in Definition 7 is replaced by multiple tapes. Typically one of those tapes contains the input, one contains the TM’s output (if and) when the TM halts, and there are one or more intermediate “work tapes” that are in essence used as scratch pads. The advantage of using this more complicated variant of TMs is that it is often easier to prove theorems for such machines than for single-tape TMs. However, there is no difference in their computational power. More precisely, one can transform any single-tape TM into an equivalent multi-tape TM (i.e., one that computes the same partial function), as well as vice-versa [5, 38, sipser2006introduction].
@@
As is standard, I fix some (prefix free, computable, invertible) encoding that maps all finite sets of finite bit strings into a single bit string. Overloading notation, I write that encoding as $\langle.\rangle$, $\langle.,.\rangle$…, depending on the number of strings in the set. Moreover, I require the input to a UTM to be encoded as $\langle k,x\rangle$ if it is emulating TM $T^{k}$ running on input string $x$. However, sometimes for clarity of presentation I will leave the angle brackets implicit, and simply write the UTM $U$ operating on input $\langle k,x\rangle$ as $U(k,x)$. In addition, as shorthand, if $x$ is a vector whose components are all bit strings, I will write the encoded version of all of its components as $\langle x\rangle$.
@@
(where “O(1)O(1)” means a term that is independent of both $x$ and $y$). Indeed, in a certain technical sense, the expected value of $K(x)$ under any distribution $P(x\in\Lambda^{*})$ equals the Shannon entropy of $P$. (See [38].)
Formally speaking, the set $\mathbb{B}$$^{*}$ is a Cantor set. A convenient probability measure on this Cantor set, sometimes called the fair-coin measure, is defined so that for any binary string $x$ the set of sequences that begin with $\sigma$ has measure $2^{-|\sigma|}$. Loosely speaking, the fair-coin measure of a prefix TM $T$ is the probability distribution over the strings in $T$’s halting set generated by IID “tossing a coin” to generate those strings, in a Bernoulli process, and then normalizing. [^23] Kraft’s inequality guarantees that since the set of strings in the halting set is a prefix-free set, the sum over all its elements of their probabilities cannot exceed $1$, and so it can be normalized. However, in general that normalization constant is uncomputable, as discussed below. Also, in many contexts we can actually assign arbitrary non-zero probabilities to the strings outside the halting set, as long as the overall distribution is still normalizable. See [38]. So any string $\sigma$ in the halting set has probability $2^{-|\sigma|}/\Omega$ under the fair-coin prior, where $\Omega$ is the normalization constant for the TM in question.
The fair-coin prior provides a simple Bayesian interpretation of Kolmogorov complexity: Under that prior, the Kolmogorov complexity of any string $\sigma$ for any prefix TM $T$ is just (the log of) the maximum over all input strings $\sigma^{\prime}$ of the joint probability of $\sigma^{\prime}$ and $\sigma$. (Strictly speaking, this result is only true up to an additive constant, given by the log of the normalization constant of the fair-coin prior for $T$.)
@@
I end this appendix with some terminological comments and definitions that will be useful in the main text. It is conventional when dealing with Turing machines to implicitly assume some invertible computable map $h(.)$ from $\mathbb{Z}$} into $\Lambda^{*}$. Given such a map $h(.)$, we can exploit it to implicitly assume an additional invertible map taking $\mathbb{Q}$ into $\Lambda^{*}$, e.g., by uniquely expressing any rational number as one product of primes, $a$, divided by a product of different primes, $b$; invertibly mapping those two products of primes into the single integer $2^{a}3^{b}$; and then evaluating $h(2^{a}3^{b})$. Using these definitions, we say that a real number $z$ is computable iff there is a recursive function $f$ mapping rational numbers to rational numbers such that for all rational-valued accuracies $\epsilon>0$, $|f(\epsilon)-z|<\epsilon$. We define computable functions from $\mathbb{Q}\rightarrow{\mathbb{R}}$ similarly.
Appendix B The recursion theorem and Rice’s theorem
@@
For any partial computable function $Q(x,y)$ there is a Turing machine with index $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$ for which $Q(x,e)$ is defined.
@@
For any TM of the form $M(y,x)$ taking two arguments, define $\llbracket{M(y,x)}\rrbracket_{x}$ as the index $e$ such that $T^{e}(x)=M(y,x)$ for all $x$, with $y$ fixed to the indicated value.
@@
(Note that $S(.)$ is a total computable function.) Using this definition, choose an index $k$ such that
$T^{k}(t,x)=Q(x,S(t))$ (44)
for all $x,t$. (At least one such index must exist since the RHS of Eq. 44 is a partial computable function.)
@@
An important special case of the theorem is where $Q(.,.)$ is a total computable function. In this case the theorem says that that there must be an $e$ such that $T^{e}(x)=Q(x,e)$ for all $x$, and therefore $T$ is also total computable. So, we can restrict to total computable functions $Q(.,)$ and total TM’s $T^{e}$, as a special case. I call this extension the “total recursion theorem” in the main text.
Kleene’s second recursion theorem has elicited some interesting commentary. For example, [49] writes that
@@
An important special case is where $G$ is restricted to a set of binary-valued partial functions. Since each of those functions is represented as a bit string, the set comprises a language. In this case Rice’s theorem concerns the decidability of languages.
Appendix C Why $\Delta t$ is not a physical variable
@@
Ultimately, the way we are getting around these problems in the framework I’ve adopted is by having $\Delta t$ just be a parameter of the evolution function, specifying how far into the future we want to simulate the evolution of the physical universe. and not a physical variable. So for example it does not need to be reproduced as the output of the computer if the cosmological universe is to simulate its own future.) A consequence of this workaround is that we need to hard-code $n^{*}$ into the initial state of the program in a way that depends on $\Delta t$. [^25] Indeed, the way the derivation of the self-simulation lemma uses the recursion theorem can be viewed as a special case of the generalized parameter-dependent form of the recursion theorem given in [49]. The derivation of the self-simulation lemma uses that generalized form for the special case where the space of parameters is single-dimensional. So in the notation of that paper, here $m=1$.
Appendix D Subtleties with the model in Section 2.6
Section 2.6 presented an example of how a portion of our actual universe could implement the purely formal model of universe $V$ given in Section 2.2. In that example the computer was implemented as a UTM, replete with tapes, etc. This might seem a particularly awkward way of modeling a portion of our cosmological universe whose dynamics is computationally universal. After all, our universe “runs in parallel”, whereas a UTM is a serial system. This means that to implement such a UTM would require copious use of energy barriers and the like, to prevent the parallel nature of Hamilton’s equations from “leaking through”.
@@
Appendix E Work in the philosophy of science literature on the simulation hypothesis
The modern iteration of the simulation hypothesis has many versions that vary in their precise details. Common to all of these versions is the idea that some portion of a physical universe, including some conscious reasoning agents that exist in that universe, might in fact be part of a simulation that is being run in a physical computer of some different, super-sophisticated reasoning agents, like some race of aliens [12, 32, 21]. [^26] A special case of this idea in which the “super-sophisticated reasoning agents” are our descendants, and so by simulating us they are simulating their ancestors, was the focus of some of the recent work on the simulation hypothesis [12]. Here I impose no such restrictions on the agents simulating us. Indeed, the physical universe in which the those reasoning agents exist might not even be one that obeys the laws of physics of our (simulated) universe. Under the simulation hypothesis, those “conscious reasoning agents” would just be variables evolving in a computer program running in that physical computer designed by that putative super-sophisticated race of aliens.
Much of philosophy of science, going back to Kant (at least), would be rendered moot if we are simulations. In particular, this would be the case for many of the flavors of “*** realism” (e.g., scientific realism), which suppose there “exists” some “real” physical truth which is not reducible to a mathematical formalism, a physical truth that instead is somehow “concrete”, and non-mathematical in its very essence — and so not just a computer algorithm. Indeed, the simulation hypothesis is consistent (in the formal logic sense) with all axiomatic systems for the laws of physics that have been offered. However, it true it would imply that there is no such “concrete”, non-mathematical reality. This establishes that it is impossible to establish the logical necessity of any of these flavors of realism, from all suggested axiomatic foundations to the laws of physics.
The central idea of the simulation hypothesis has been extended in an obvious manner, simply by noting that the aliens that simulate our universe might themselves be a simulation in the computer of some even more sophisticated species, and so on and so on, in a sequence of ever-more sophisticated aliens. Similarly, going the other direction, in the not too distant future it is conceivable that we would be able to produce our own simulation of a universe running in some future computer that we will create, a simulation complete with variables that constitute “conscious, reasoning agents”. Indeed, we might produce such a simulation in which the reasoning agents can produce their own simulated universe in turn, etc. If this were the case, then in the near future there might be a sequence of species’, each one with a computer running a simulation that produces the species just below it in the sequence, with us somewhere in the “middle” of that sequence.
The literature of the last few decades on the simulation hypothesis has focused almost exclusively on whether we, in our universe are such a simulation. This question is answered rather trivially if we adopt the view of ontic structural realism [44, 13], especially if that view is formalized in terms of Tegmark’s level IV multiverse [63, 64]: yes, in some universes we are a simulation, and no, in some other universes we are not. (See also [17, 54].)
These investigations of relative probabilities all beg (many) questions about what such probabilities might actually mean, formally. (Just as there are such questions concerning the idea of ascribing probabilities to the universes in a level IV multiverse in general.) How does one ascribe a measure, obeying the Kolmogorov axioms based on some associated sigma algebra, to a collection of events each of which is a universe? Note that it is not even clear that this collection would be a set, rather than a proper class of some sort. Directly reflecting such problems, one could not experimentally assess any proposed value of such a probability, e.g., with a proper scoring rule. [^27] One might try to use a Bayesian “degree of belief” interpretation of probability to circumvent this issue. However, without any decision that would be made based on the probability assigned, and associated loss function, one could not apply Bayesian decision theory. So it is hard to see how “degree of belief” is meaningful in this case. See [67] for some related arguments.
In particular, many papers considering such probabilities assume, implicitly or otherwise, that there is some way to assign a uniform probability to all universes. However, it is proven in Section 2.4 that this is mathematically impossible in sets with the cardinality of the kind of universes considered in these other papers. One might imagine that rather than a uniform probability distribution, one could assign some sort of Cantor (“fair coin”) measure to the set of possible universes, as is done for example in algorithmic information theory. However, it is proven in Section 7 that it is impossible to assign a Cantor measure to a set of universes obeying a natural set of restrictions.
Yet others have considered issues concerning the simulation hypothesis that might best be characterized as “philosophical” in nature [33, 22, 27, 39], This work considers potential normative implications if the simulation hypothesis holds, or even such implications that follow from the mere possibility that the simulation hypothesis holds.
Appendix F Semi-formal work that has been done on the CS theoretic aspects of the simulation hypothesis
The only work that has been done that is even tangentially related to the CS aspects of the simulation hypothesis was a set of semi-formal results presented in [70]. Some of those semi-formal results define “simulation” in terms of the relationship between languages at different levels of the time, polynomial, or exponential hierarchies of CS theory, and some of them define “simulation” in terms of the relationship between languages at different levels the arithmetic or analytic hierarchies of logic, or similar constructions. (See [23, 5, 31] for background on such hierarchies.) Other results in [70] define “simulation” in terms of the relationship between computational machines with different Turing degrees [62, 60]. Yet others consider the implications of Gödel’s second incompleteness theorem, defining simulation in terms of languages (in the logic theory sense) with nested sets of axioms.
None of these results in [70] consider what it means for a computational machine to simulate a physical universe, per se. There is no concern for “coupling” the mathematics of CS theory to the laws of physics of our universe. No work has been done to date concerning the simulation hypothesis that focuses on the dynamical system of an evolving physical universe, like ours.